OpenStudy (anonymous):

By writing x as 4+(x-4) and using the Maclaurin expansion for ln(1+t), or otherwise, find the first four non zero terms in the Taylor series expansion for ln(5x) about x = 4.

OpenStudy (jamesj):

ln(5x) = ln 5 + ln x = ln 5 + ln(4 + (x-4)) = ln 5 + ln 4 + ln(1 + (x-4)/4) = ln(20) + ln(1 + u), where u = (x-4)/4 Now use the series expansion for ln(1+u)